Solve the system of linear equations below.
x + y = 4
2x + 3y = 0

Answer by checkley79(3341)

x+y=4 MULTIPLY BY -2 & ADD.
2x+3y=0
-2X-2Y=-8
-------------------------
Y=-8 ANS.
2X+3*-8=0
2X-24=0
2X=24
X=24/2
X=12 ANS.
PROOF:
12-8=4
4=4